# Accelerating Training Speed of Tiny Recursive Models via Curriculum-Guided Adaptive Recursion

### v1

@@ VI Conclusion
We have presented CGAR (Curriculum-Guided Adaptive Recursion), the first training methodology to apply curriculum learning directly to architectural recursion depth rather than traditional data ordering. Through progressive depth curriculum and hierarchical supervision weighting, CGAR fundamentally rethinks how recursive reasoning models learn. Our experimental evaluation on Sudoku-Extreme demonstrates $1.71\times$ training speedup ($6.38$ hours vs $10.93$ hours, 42% cost reduction) while maintaining competitive accuracy ($86.02\%$ on 423,168423{,}168 test puzzles). Systematic ablation studies reveal that progressive depth curriculum is the dominant component achieving $2.26\times$ speedup with comparable 85.47% accuracy versus baseline’s 85.14%, demonstrating a rare Pareto improvement where architectural curriculum simultaneously enhances training efficiency and maintains solution quality. Hierarchical supervision weighting provides complementary learning efficiency gains ($1.61\times$ speedup through 40% gradient variance reduction). Notably, the curriculum-only configuration achieves both faster training and comparable accuracy, challenging the conventional wisdom that efficiency and quality are opposing objectives.
The key innovation is conceptually elegant yet powerful: treating architectural depth $\mathcal{D}_{\text{eff}}(n,T)$ as a curriculum-driven training hyperparameter rather than a fixed constant. By applying curriculum learning to recursion depth rather than data ordering, CGAR enables models to start with shallow, computationally efficient iterations for fundamental pattern learning, then progressively increase recursion depth to handle complex logical dependencies as optimization proceeds. This paradigm shift mirrors human learning principles and aligns with information-theoretic observations that early reasoning steps carry disproportionate gradient information. Our findings demonstrate that curriculum-trained models achieve superior inference efficiency with perfect halting decisions (100% Q-halt accuracy) and $11\%$ fewer reasoning steps, suggesting that progressive depth curriculum teaches not just what to predict, but when to engage deep reasoning versus shallow inference. CGAR demonstrates that principled curriculum on architectural depth, rather than traditional data-centric approaches, can fundamentally accelerate recursive reasoning model training while maintaining or improving model quality.
By reducing training costs by $42\%$ (\$21.86 $\to$ \$12.76 per run), CGAR makes recursive reasoning models more accessible to researchers with limited compute budgets, enabling faster iteration cycles and broader experimentation in neurosymbolic AI, program synthesis, and interpretable reasoning systems. The principles demonstrated here curriculum on architecture, recursion-aware supervision may apply broadly to training efficient foundation models with adaptive computation. We hope this work inspires further research into architectural curriculum learning and training methodologies that improve both efficiency and model quality for next-generation reasoning systems.
@@ VI-A Limitations and Future Work
While our results on Sudoku-Extreme are promising, broader validation across diverse reasoning tasks would strengthen the generalizability claims. Due to computational budget constraints, we focused on a single task domain; future work should evaluate CGAR on ARC-AGI benchmarks, mathematical reasoning datasets (GSM8K, MATH), and other constraint satisfaction problems to establish task-agnostic effectiveness. The curriculum schedule thresholds $(\tau_{1},\tau_{2})=(0.3,0.6)$ and depths $(n,T)\in\{(2,1),(4,2),(6,3)\}$ were manually tuned on Sudoku-Extreme; automated curriculum optimization through meta-learning or validation-based transitions could improve adaptability across tasks with different complexity profiles.
Future research directions include sample-adaptive depth allocation, where recursion depth adjusts per-example based on puzzle difficulty rather than uniform epoch-based scheduling, potentially reducing computational waste on easy instances. Theoretical analysis characterizing optimal curriculum schedules for different loss landscape geometries would provide principled design guidelines beyond empirical tuning. The subadditive interaction between PDC and HSW ($1.71\times<2.26\times 1.61$) warrants deeper investigation into gradient flow dynamics during depth transitions to improve component synergy. Finally, extending CGAR principles to large-scale pretraining scenarios could demonstrate whether architectural curriculum applies beyond task-specific training, potentially reducing multi-thousand GPU-hour costs for foundation models while improving optimization trajectories.

### v2

@@ VI Conclusion
This paper has presented CGAR (Curriculum-Guided Adaptive Recursion), a training methodology that applies curriculum learning to architectural recursion depth rather than data ordering. CGAR consists of two components: Progressive Depth Curriculum dynamically adjusts recursion parameters $(n,T)$ from shallow to deep configurations during training, while Hierarchical Supervision Weighting applies exponentially decaying importance $w_{t}=\lambda^{t-1}$ to supervision steps based on observed gradient magnitude decay in recursive architectures. Experimental evaluation on Sudoku-Extreme with 423,168 test puzzles demonstrates $1.71\times$ training speedup (10.93 to 6.38 hours) with 0.63% accuracy reduction (86.65% to 86.02%). Systematic ablation studies reveal Progressive Depth Curriculum achieves $2.26\times$ speedup with comparable accuracy (85.47% vs 85.14% baseline), while Hierarchical Supervision Weighting provides $1.61\times$ speedup through 40% gradient variance reduction.
The approach treats architectural depth $\mathcal{D}_{\text{eff}}(n,T)$ as a curriculum-scheduled training parameter rather than a fixed constant. This enables computational savings during training while preventing early-stage overfitting. CGAR-trained models demonstrate improved inference efficiency with 100% halting accuracy and 11% fewer reasoning steps compared to baseline. The 42% training cost reduction improves accessibility for resource-constrained research environments. By reducing training time from 10.93 to 6.38 hours on standard hardware, CGAR makes recursive reasoning models more practical for broader adoption in neurosymbolic AI, program synthesis, and interpretable reasoning systems.
@@ VI-A Limitations and Future Work
While our results on Sudoku-Extreme are promising, broader validation across diverse reasoning tasks would strengthen generalizability claims. Due to computational constraints, we focused on a single task domain. Future work should evaluate CGAR on ARC-AGI benchmarks and other constraint satisfaction problems to establish task-agnostic effectiveness. The curriculum schedule thresholds $(\tau_{1},\tau_{2})=(0.3,0.6)$ and depths $(n,T)\in\{(2,1),(4,2),(6,3)\}$ were manually tuned on Sudoku-Extreme; automated curriculum optimization through meta-learning or validation-based transitions could improve adaptability across tasks with different complexity profiles.
Future research directions include sample-adaptive depth allocation, where recursion depth adjusts per-example based on instance difficulty rather than uniform epoch-based scheduling. Theoretical analysis characterizing optimal curriculum schedules for different loss landscape geometries would provide principled design guidelines beyond empirical tuning. The subadditive interaction between PDC and HSW ($1.71\times<2.26\times 1.61$) warrants investigation into gradient flow dynamics during depth transitions. Finally, extending CGAR principles to large-scale pretraining scenarios could demonstrate whether architectural curriculum applies beyond task-specific training, potentially reducing multi-thousand GPU-hour costs for foundation models.